# Convolutional Kolmogorov–Arnold Networks

### v2

@@ Abstract
In this paper, we introduce Convolutional Kolmogorov-Arnold Networks (Convolutional KANs), an innovative alternative to the standard Convolutional Neural Networks (CNNs) that have revolutionized the field of computer vision. By integrating the learneable non-linear activation functions presented in Kolmogorov-Arnold Networks (KANs) into convolutions, we propose a new layer. Throughout the paper, we empirically validate the performance of Convolutional KANs against traditional architectures across Fashion-MNIST dataset, finding that, in some cases, this new approach maintains a similar level of accuracy while using half the number of parameters. KAN Convolutions seem to learn more per kernel, which opens up a new horizon of possibilities in deep learning for computer vision.
@@ 1 Introduction
The field of deep learning is constantly changing, the fast improvement of architectures has helped the advancement of computer vision in tasks involving complex spatial data. Convolutional Neural Networks proposed by LeCun et al.[8] are widely used due to their ability to handle high-dimensional data arrays such as images. Normally, these networks rely on linear transformations followed by an optional activation function in their convolutional layers to understand spatial relationships, which significantly reduced the number of parameters to capture complex patterns in images.
In recent years, there has been an increase in the integration of advanced mathematical theories into deep learning architectures which have helped neural networks in handling complex data structures. Kolmogorov-Arnold Networks (KANs) [10] are a promising alternative to Multi-Layer Perceptrons (MLPs)[6] that use the Kolmogorov-Arnold theorem to integrate splines which is a key component of their architecture.
In light of these advancements, this paper explores the adaptation of KANs to convolutional layers, a common element in many CNN architectures used in computer vision. Traditional CNNs utilize fixed activation functions and linear transformations which, while effective, can benefit from the flexibility and reduced parametric complexity offered by KANs. By employing spline-based convolutional layers, as proposed in SplineCNN by M. Fey and J. E. Lenssen et al. [4], networks can capture non-linear relationships more effectively.
Throughout this paper, we begin with a high-level overview of the KAN architecture to set the stage for a comprehensive mathematical treatment of Convolutional KANs. We will provide a detailed examination of different Convolutional KANs architectures and benchmark their performance against traditional models, focusing on parameter efficiency within the MNIST and Fashion-MNIST datasets. Our hypothesis posits that Convolutional KANs, by leveraging spline-based layers, will require fewer parameters while achieving accuracy levels competitive with established benchmarks, potentially setting a new standard in neural network architectures for image-related tasks [3]. For further exploration and practical application, the code for this layer and all the experiments is available at our GitHub repository: GitHub/Convolutional-KANs.
@@ 3.1 Architecture
In more detail, the Kolmogorov-Arnold representation theorem posits that a multivariate function $f(x_{1},\ldots,x_{n})$ can be expressed as:
@@ 4.1 KAN Convolutions
In Computer Vision, convolutions are normally used interchangeably with a mathematical operation used in Convolutional Neural-Networks. The operation consists in passing a kernel, or filter, across the input and calculating the dot product at each position. In KAN Convolutions the main idea is to propose an alternative implementation of this mathematical operation utilizing the approach of Kolmogorov-Arnold Networks. The main difference between KAN Convolutions and the convolutions used in CNNs lies in the kernel. In CNNs it is made of weights whereas in Convolutional KANs, each element of the kernel, $\phi$, is a learnable non-linear functions that utilizes B-Splines. Formally, each element is defined as:
@@ 5 Experiments
We conducted a hyperparameter search using grid search, testing 8 different combinations for each model. For KAN models, we performed different grid searches for KAN grid sizes of 10 and 20 (the grid size for the B-Spline) , as parameter count increases significantly with grid size and we wanted to compare how the accuracy changes when grid size changes. Another possible hiperparameter is the B-Spline degree. We decided to not tune this hiperparameter and chose it as degree 3, as KAN authors [10] suggest as default and because of the wide usage of cubic splines. It is important to define the degree sufficiently small to not overfit.
Because of the long training times of these models, we opted not to do a K-Fold Cross Validation, but to split the training set into train and valid sets and tune hiperparameters with those sets. Once we find the best hiperparameter sets with the previous datasets, we join train and valid together, train a model with train and valid sets with the optimal hiperparameters and finally report the results using test set.
@@ 5.1 Loss
In our early experiments we found out that the trained models performed better and trained faster wihtout the regularization terms, so in our experiments we ended defining $\lambda=0$.
@@ 5.2 Hardware
For conducting the hiperparameter tuning for all the models, we used a ’g2-standard-4’ Google Cloud instance, which has 4 VCPUS, 2 cores, 16 RAM and a NVIDIA L4 GPU.
@@ 6 Results
This section displays an analysis of the performance of the different proposed models in the previously described experiments. The following Figure LABEL:fig:Accuracy_datasets shows the final accuracies of the proposed models. To obtain the hiperparameters and final models, the training time was of almost 2 and a half days.
Table LABEL:tab:model_performances_Fashion_MNIST presents a comparison of accuracy, precision, recall, F1 score, parameter count, and training time per epoch for the proposed models tested on the Fashion MNIST Dataset.
@@
Results of the table, can be seen in Figure 4 which illustrates the relationship between parameter count and accuracy for these models, providing a graphical representation of the data.
@@
Following the results analysis, Figure 4 and Table LABEL:tab:model_performances_Fashion_MNIST illustrate the relationship between parameter counts versus accuracy for the described models but applied to the Fashion MNIST dataset. In the experiments with this dataset, we find that:
@@
Although KANC MLP and CNN of equal sizes have the same architecture and only changes the type of convolution, KAN convolutions provide a considerable ammount of parameters. Ignoring the difference of parameters, as said before, within the same architechture, both Small and Medium KANC MLPs outperform CNNs. But, for example, KANC MLP (Small) is bigger than CNN (Medium) and with slighly less accuracy ($88.34\%$ vs $88.15\%$).
@@
In the case where we use a KAN Network after the convolutions, using KAN convolutions (KKAN Small) gets $87.67\%$ accuracy while using classic convolutions, Conv & KAN (Small) gets $88.01\%$ and Conv & KAN (Medium) gets $87.92\%$ Conv & KAN is 6 or 7 times faster to train (depending on the grid size), but KKAN Small achieves slightly better accuracy than Conv & KAN (Medium) (for any grid sizes) and with almost half parameters (when Conv KAN (Medium) has grid size 20 and KKAN grid size 10). Needless to say, KKAN (Medium), which only increases the ammount of convolutions, suprasses this accuracy with grid size 20, which achieves $88.56\%$, but parameter count increases from 38000 to 74875 (Because of the increased number of convolutions and KAN Neurons).
@@
In addition, we can verify that grid size can change abrutly the model accuracy. For achieving the best performance this parameter should be tuned for every KAN layer, although this process demands significantly greater computational resources. In the cases of KKAN Small, KANC MLP Medium and Big and both Conv & KANs, no accuracy was gained by increasing grid size from 10 to 20, and in some instances, accuracy was even reduced. However, in models such as CNN Small and KKAN Medium, accuracy improved by over $0.5\%$, which represents a noticeable gain given the minimal accuracy differences across these models.
@@ 7 Conclusions
In this paper we proposed a new way to adapt the idea of learning splines proposed in Kolmogorov-Arnold Networks to convolutional layers used widely in Computer Vision. We implemented a KAN Convolutional Layer that uses a kernel made of learnable non-linear functions that use B-Splines. We have found out that with equal architectures, KAN Convolutions seem to "learn more", showing better accuracy in the analogous models except in the "Big" ones, where the difference was that we added more fully connected layers, which may be doing the learning instead of the convolutions.
When using KANs after the flatten, KAN Convolutions achieve better accuracy than we can classic convolutions even using half the parameters. This is seen in the comparisons between KKAN and Normal Conv & KAN.
When trying an MLP, KAN convolutions achieve higher accuracy in the small models, but when having a 2 layer MLPs, the classic CNN wins by $0.41\%$ with $\sim 26.62$k parameters. While KAN Convolutions seem to learn more per kernel, we have to consider that each KAN kernel has much more parameters. So comparing the same architectures gives an advantage in terms of expressibility to KAN Convolutions.
A key factor here is that KANs seem to maintain accuracy with lower parameter count, KANC MLP (Medium) achieves $88.99\%$ with almost $15k$ parameters, but the training time is almost 10 times slower with the current implementations of KANs and its derivatives.
Additionally, MLPs show a better performance than KANs for image classification tasks. In the "Small" and "Medium" cases, when using KAN Convolutions, using MLPs after the flatten gives both better accuracy and smaller parameters counts than using KANs. As expected, KKANs achieve better accuracy than its equal CNNs, but the parameter count difference is too big (22k vs 1.5k and 38k vs 3k). Since the parameters of KAN layers grow much faster than those of MLPs, with no apparent gain in accuracy, KANs might not be suitable as dense layers, especially because CNNs require a large number of neurons after the flattening layer.
Based on the original KAN paper [10] proposal that KANs are more interpretable, we have tried to find a way to interpret this new type of convolutions, but at the moment, we have not found any clear way to visualize the B-Splines learned in each pixel of the convolutions. Given that we are working on images, it seems that the classic approach of visualizing what the filter does to the image seems the more ’human’ way to get a sense of what is being learned, but because of the high dimensionality of a KAN kernel input-output relationship, at the moment we haven’t come up with any interpretation.
@@ 8 Limitations and future work
The current implementation of KANs with B-Splines is considerably slow due to its impossibility of being GPU parallelizable, making it very difficult to apply KANs in real world problems. Many authors are working on solutions to this by replacing B-Splines by other function approximators, such as Radial Basis Function LABEL:radialbasis, which are GPU parallelizable and open up a wide range of possibilities with KANs.

### v3

# Convolutional Kolmogorov-Arnold Networks

@@ Abstract
In this paper, we present Convolutional Kolmogorov-Arnold Networks, a novel architecture that integrates the learnable spline-based activation functions of Kolmogorov-Arnold Networks (KANs) into convolutional layers. By replacing traditional fixed-weight kernels with learnable non-linear functions, Convolutional KANs offer a significant improvement in parameter efficiency and expressive power over standard Convolutional Neural Networks (CNNs). We empirically evaluate Convolutional KANs on the Fashion-MNIST dataset, demonstrating competitive accuracy with up to $50\%$ fewer parameters compared to baseline classic convolutions. This suggests that the KAN Convolution can effectively capture complex spatial relationships with fewer resources, offering a promising alternative for parameter-efficient deep learning models.
@@ 1 Introduction
The field of deep learning is constantly changing, the fast improvement of architectures has helped the advancement of computer vision in tasks involving complex spatial data. Convolutional Neural Networks proposed by LeCun et al.[8] are widely used due to their ability to handle high dimensional data arrays such as images. Normally, these networks rely on linear transformations followed by an optional activation function in their convolutional layers to understand spatial relationships, which significantly reduced the number of parameters to capture complex patterns in images.
Recent years have seen an increase in the integration of advanced mathematical theories into deep learning architectures which have helped neural networks in handling complex data structures. Kolmogorov-Arnold Networks (KANs) [10] are a promising alternative to Multi-Layer Perceptrons (MLPs)[6] that use the Kolmogorov-Arnold theorem to integrate splines which is a key component of their architecture.
In light of these advancements, this paper explores the adaptation of KANs to convolutional layers, a common element in many CNN architectures used in computer vision. Traditional CNNs utilize fixed activation functions and linear transformations which, while effective, can benefit from the flexibility offered by KANs. By employing spline-based convolutional layers, as proposed in SplineCNN by M. Fey and J. E. Lenssen et al. [4], networks can capture non-linear relationships more effectively.
Throughout this paper, we begin with a high-level overview of the KAN architecture to set the stage for a comprehensive mathematical treatment of Convolutional KANs. We will provide a detailed examination of different Convolutional KANs architectures and benchmark their performance against traditional models, focusing on parameter efficiency within the Fashion-MNIST dataset. Our hypothesis posits that Convolutional KANs, by leveraging spline-based layers, will require fewer parameters while achieving accuracy levels competitive with established benchmarks, potentially setting a new standard in neural network architectures for image-related tasks [3]. For further exploration and practical application, the code for this layer and all the experiments is available at our GitHub repository: GitHub/Convolutional-KANs.
@@ 3.1 Architecture
In more detail, the Kolmogorov-Arnold representation theorem posits that a continous multivariate function $f(x_{1},\ldots,x_{n})$ can be expressed as:
@@ 4.1 KAN Convolutions
In computer vision, convolutions are a foundational operation in Convolutional Neural Networks (CNNs), where a kernel or filter slides across the input, computing dot products at each position to extract spatial features. Unlike traditional CNNs, where kernels consist of fixed weights, KAN Convolutions use kernels composed of learnable non-linear functions, denoted by $\phi$, which are parameterized using B-splines. This approach allows each kernel element to adapt dynamically during training, enabling greater flexibility and expressiveness. Formally, each element of the kernel is defined as:
@@
4.2 KAN filters theoretical advantages
The early application of different nonlinearities to each pixel makes KAN kernels able to do things that classic kernels cannot. Because each pixel can be transformed nonlinearly before being summed, the resulting filter can encode complex, context-dependent behaviors in a single pass. By contrast, classic linear kernels merely combine pixel intensities through fixed, constant weights. While multiple linear filters can be stacked or summed, they still cannot implement conditional or piecewise logic at the pixel level without separate processing steps. KAN filters, however, can carry out thresholding, adaptive smoothing, and edge detection all at once.
Let’s take a look at some of these examples to illustrate the unique capabilities of KAN filters:
4.2.1 Threshold + Convolution
Suppose we want to:
  1. 1.
    Apply a threshold on specific pixels (e.g., the center pixel).
  2. 2.
    Linearly weight each pixel’s intensity. This would also apply to other desired operations like Sobel or Laplacian filters.
A single classical filter cannot perform a threshold in the same pass (that requires a global post-sum activation or a separate step). In KAN form, however, we define each $\phi_{i,j}$ as:
$\phi_{i,j}(x)\;=\;\begin{cases}w_{i,j}\,x,&\text{if $(i,j)$ is not the center},\\[6.0pt] \max(0,x-1)&\text{if $(i,j)$ is center}.\end{cases}$
Here, the second function looks like a displaced ReLU, which doesn’t let any positive value between 0 and 1 pass, and then progressively activates. To clarify, this would not be exactly the learned function because we can only learn smooth functions, but it can be sufficiently close in practice.
Summing $\phi_{i,j}(x_{i,j})$ over all $i,j$ yields a final image where we both apply a linear weighting and smooth thresholding to the center pixel in one pass.
4.2.2 Piecewise Brightness-Based Boost
Although $\phi_{i,j}$ cannot see neighboring pixels, it can still apply a different multiplier depending on whether its own intensity is above or below a threshold. For instance:
$\phi_{i,j}(x)\;=\;\begin{cases}\alpha_{\text{bright}}\,\cdot\,x,&\text{if }x\geq T,\\[6.0pt] \alpha_{\text{dark}}\,\cdot\,x,&\text{otherwise}.\end{cases}$
Here:
  • $\alpha_{\text{bright}}$ and $\alpha_{\text{dark}}$ are different gains for bright vs. dark pixel intensities.
  • If $x_{i,j}\geq T$, we “boost” it more ($\alpha_{\text{bright}}$), else we “boost” less ($\alpha_{\text{dark}}$).
Summing $\phi_{i,j}(x_{i,j})$ over $i,j$ yields an output that conditionally scales each pixel based on its own brightness—something no single linear filter can replicate. For instance, if $\alpha_{\text{bright}}>\alpha_{\text{dark}}$, bright regions get emphasized more strongly than dark regions in one pass.
4.2.3 Key takeaways
KAN filters, even with only per-pixel access, allow us to mix linear and nonlinear operations (thresholding, piecewise gains, small saturations, etc.) all in a single pass—whereas a single classic filter must remain purely linear, and any nonlinear steps would typically happen after that pass and without the possibility of letting gradient descent learn the activation function. Also, if we would want to achieve the same results as the KAN Convolution, we would need to apply pixel operations before the convolution, which is not typically done in CNNs.
@@ 5 Experiments
We conducted a hyperparameter search using grid search, testing 8 different combinations for each model. For KAN models, we performed different grid searches for KAN grid sizes of 10 and 20 (the grid size for the B-Spline), as parameter count increases significantly with grid size and we wanted to compare how the accuracy changes when grid size changes. Another possible hyperparameter is the B-Spline degree. We decided to not tune this hyperparameter and chose it as degree 3, as KAN authors [10] suggest as default and because of the wide usage of cubic splines. It is important to define the degree sufficiently small to not overfit.
Because of the long training times of these models, we opted not to do a K-Fold Cross Validation, but to split the training set into train and valid sets and tune hyperparameters with those sets. Once we find the best hyperparameter sets with the previous datasets, we join train and valid together, train a model with train and valid sets with the optimal hyperparameters and finally report the results using test set.
@@ 5.1 Loss
In our early experiments we found out that the trained models performed better and trained faster without the regularization terms, so in our experiments we ended defining $\lambda=0$.
@@ 5.2 Hardware
For conducting the hyperparameter tuning for all the models, we used a ’g2-standard-4’ Google Cloud instance, which has 4 VCPUS, 2 cores, 16 RAM and a NVIDIA L4 GPU.
@@ 6 Results
This section displays an analysis of the performance of the different proposed models in the previously described experiments. To obtain the hyperparameters and final models, the training time was of almost 2 and a half days. Table 2 presents a comparison of accuracy, precision, recall, F1 score, parameter count, and training time per epoch for the proposed models tested on the Fashion MNIST Dataset.
@@
Results of Table 2 , can also be visualized in Figure 4 which illustrates the relationship between parameter count and accuracy achieved by these models, providing a graphical representation of the results.
@@
Following the results analysis, Figure 4 and Table 2 illustrate the relationship between parameter counts versus accuracy for the described models but applied to the Fashion MNIST dataset. In the experiments with this dataset, we find that:
@@
Although KANC MLP and CNN of equal sizes have the same architecture and only changes the type of convolution, KAN convolutions provide a considerable amount of parameters. Ignoring the difference of parameters, as said before, within the same architecture, both Small and Medium KANC MLPs outperform CNNs. But, for example, KANC MLP Small has more trainable parameters than CNN Medium and had slightly less accuracy ($88.15\%$ vs $88.34\%$). While it has more parameters, it has less convolution kernels, so in this case, it seems that the added expressibility of KAN Convolutions does not compensate the number of different filters.
@@
In the case where we use a KAN Network after the convolutions, using KAN convolutions (KKAN Small) gets $87.67\%$ accuracy, while using classic convolutions (Conv & KAN Small) gets $88.01\%$. Conv & KAN is between 6 and 7 times faster to train (depending on the grid size), but KKAN Small achieves slightly better accuracy than Conv & KAN Medium ($87.92\%$) and with almost half parameters (when Conv KAN Medium has grid size 20 and KKAN grid size 10). Needless to say, KKAN Medium, which only increases the amount of convolutions, surpasses this accuracy with grid size 20, which achieves $88.56\%$, but parameter count increases from 38000 to 74875 (Because of the increased number of convolutions and KAN Neurons).
@@
In addition, we can verify that grid size can change abruptly the model accuracy. For achieving the best performance this parameter should be tuned for every KAN layer, although this process demands significantly greater computational resources. In the cases of KKAN Small, KANC MLP Medium and Big and both Conv & KANs, no accuracy was gained by increasing grid size from 10 to 20, and in some instances, accuracy was even reduced. However, in models such as KANC MLP Small and KKAN Medium, accuracy improved by over $0.5\%$, which represents a noticeable gain given the minimal accuracy differences across these models.
@@ 7 Conclusions
In this paper we proposed a new way to adapt the idea of learning splines proposed in Kolmogorov-Arnold Networks to convolutional layers used widely in Computer Vision. We implemented a KAN Convolutional Layer that uses a kernel made of learnable non-linear functions using B-Splines. We have found out that with equal architectures, KAN Convolutions seem to “learn more", showing better accuracy in the analogous models except in the “Big" ones, where the difference was that we added more fully connected layers, which may be doing the learning instead of the convolutions.
When using KANs after the flatten, KAN Convolutions achieve better accuracy than classic convolutions, even using half the parameters. This is seen in the comparisons between KKAN and Normal Conv & KAN.
When trying a MLP, KAN convolutions achieve higher accuracy in the small models, but when having a 2 layer MLPs, the classic CNNs win by $0.41\%$ with $\sim 26.62$k parameters. While KAN Convolutions seem to learn more per kernel, we have to consider that each KAN kernel has much more parameters. So comparing the same architectures gives an advantage in terms of expressibility to KAN Convolutions.
A key factor found is that KANs seem to maintain accuracy with lower parameter count, KANC MLP (Medium) achieves $88.99\%$ with approximately $9k$ parameters, in comparison to CNN (Big), which reaches $89.44\%$ with $26k$ parameters. But the training time is almost 8 times slower with the current implementations of KANs and its derivatives.
Additionally, MLPs show a better performance than KANs for image classification tasks. In the “Small" and “Medium" cases, when using KAN Convolutions, using MLPs after the flatten gives both better accuracy and smaller parameter counts than using KANs. As expected, KKANs achieve slightly better accuracy than its equal CNNs, but the parameter count difference is too high (22k vs 1.5k and 38k vs 3k), without a corresponding improvement in accuracy ($0.91\%$ increase in accuracy for the Small case and $0.22\%$ for the Medium case). Since the parameters of KAN layers grow much faster than those of MLPs, with no significant gain in accuracy, KANs might not be suitable as dense layers, especially because CNNs require a large number of neurons after the flattening layer.
Based on the original KAN paper [10] proposal that KANs are more interpretable, we have tried to find a way to interpret this new type of convolutions, but at the moment, we have not found any clear way to visualize the B-Splines learned in each pixel of the convolutions besides that KAN Convolutions are able by constructions to learn more filters in one kernel. Given that we are working on images, it seems that the classic approach of visualizing what the filter does to the image seems the most ’human’ way to get a sense of what is being learned.
@@ 8 Limitations and future work
The current implementation of KANs with B-Splines is considerably slow due to its impossibility of being GPU parallelizable, making it very difficult to apply KANs in real world problems. Many authors are working on solutions to this by replacing B-Splines by other function approximators, such as Radial Basis Function [9], which are GPU parallelizable and open up a wide range of possibilities with KANs.